OpenStudy (anonymous):

A block of wood has a length of 4 cm, a width of 5 cm, and a height of 10 cm. What is the volume of the wood? 20 cm³ 19 cm³ 50 cm³ 200 cm³

OpenStudy (anonymous):

@ZairaHernandez @timo619 @Tyler.football.player

OpenStudy (anonymous):

A block, meaning cube, multiply all of the dimensions, V = l*h*w

OpenStudy (anonymous):

Notice that volume is just area*height

OpenStudy (anonymous):

\[V = l \times w \times h\]

OpenStudy (anonymous):

whe answer all i need

OpenStudy (anonymous):

That's not how it works here, you have the information.

OpenStudy (anonymous):

its in the question

OpenStudy (anonymous):

I have gave you the required formula, all you have to do is now just plug and chug...

OpenStudy (anonymous):

omg ur are no help bye

OpenStudy (anonymous):

Well I can always suspend you if you're only here for the answers.

OpenStudy (anonymous):

I suggest you read the CoC, which you signed when you joined again. http://openstudy.com/code-of-conduct